OpenStudy (anonymous):

The slope of a line is 1, and the y-intercept is -1. What is the equation of the line written in slope-intercept form? y = x - 1 y = 1 - x y = -x - 1

OpenStudy (caozeyuan):

y=kx+b, where k is the slope and b is the y-intercept

OpenStudy (anonymous):

Are you familiar with slope-intercept, or y = mx + b ?

OpenStudy (caozeyuan):

I use k for some reason, but m works just fine

OpenStudy (anonymous):

Yes, but would it just be y= x -1?

OpenStudy (caozeyuan):

which is the first

OpenStudy (mayankdevnani):

try to explain more @caozeyuan

OpenStudy (caozeyuan):

sorry eng is not my native language so I am knidn of limited

OpenStudy (mayankdevnani):

@tdakota do you satisfy with explanations ?